Determine the average value of f(x,y) = ex+y over the region R=[0,2]X[0,2].


1
Expert's answer
2021-05-07T14:22:52-0400

Let "f(x,y)=e^{x+y}" and "R=[0,2]\times [0,2]" .

The average value of "f" over "R" is "f_\text{avg}=\frac{1}{\text{Area}_R}\iint _R f(x,y)dA" .

"\text{Area}_R=2^2=4"

"\iint _R f(x,y)dA=\int\limits_0^2\int\limits_0^2e^{x+y}dxdy=\int\limits_0^2 \int\limits_0^2 e^{x}\cdot e^ydx dy=
\int\limits_0^2e^{x}dx \cdot \int\limits_0^2e^ydy=e^x\big|^2_0\ \cdot \ e^y\big|_0^2=(e^2-1)^2"


Answer: "f_{\text{avg}}=\frac{(e^2-1)^2}{4}" .